OpenStudy (anonymous):

find the numerical value for W such that L hopital rule applies to the following limit then compute value L of that limit lim x-->3 4x^(2)-10x +W/x-3 =L

OpenStudy (tkhunny):

If you mean \(\dfrac{4x^{2}-10x+W}{x-3}\)? That is NOT what you have written.

OpenStudy (tkhunny):

\(\lim\limits_{x\rightarrow3} (x-3) = 0\) Find \(W\) so that: \(\lim\limits_{x\rightarrow3} (4x^{2} - 10x + W) = 0\)

OpenStudy (anonymous):

so will W be 6

OpenStudy (anonymous):

-6

OpenStudy (raden):

agreed ^^

OpenStudy (tkhunny):

That will do, if I have translated the expression correctly.

OpenStudy (anonymous):

but -6 is wrong

OpenStudy (anonymous):

what happened to L? This function is equal to L.

OpenStudy (tkhunny):

Did you answer the question? Is asks for the limit, not just the value of W.

OpenStudy (tkhunny):

Ever graph a Rational Function?

OpenStudy (anonymous):

how do you find L?

OpenStudy (tkhunny):

You didn't answer my question. I'm trying to answer yours.

OpenStudy (anonymous):

i got it it. i factored the top function and which will cancel x-3 in the denominator. thanks!

OpenStudy (tkhunny):

Good call. Does the problem statement DEMAND that you use l'Hopital's Rule?